# Fully Dynamic Euclidean $k$-Means

### v2

@@ 4.3 Approximate Assignment Structure
However, there is no finite ratio with small recourse for this task, since one can keep on inserting/deleting a point of $S$ to change the assignment drastically. To bypass this barrier, instead of maintaining the full assignment $\sigma:S\to X$, our idea is to hash the dataset into buckets $\varphi(X)$ and explicitly maintain only an intermediate assignment $\sigma^{\prime}$ between bucket (ID) and $S$. The final assignment $S\to X$ is implicitly defined by assigning all points in the same bucket to the same bucket partner in the intermediate assignment.
@@
Lemma 4.7 implies key lemmas that we use for restricted $k$-Means and augmented $k$-Means (see Section 4.4 for the definitions of these sub-tasks). Similar to Lemma 4.7, these results maintain dynamic $X$ and $S$, work against an adaptive adversary and succeeds with high probability.
Application: Cluster Weight Estimation. An immediate application is to compute the injectivity $w_{s}$ of points $s\in S$ in the assignment $\sigma$ (which is $\mathrm{poly}(\Gamma)$-approximate by Property 1 of Lemma 4.7), via maintaining $|\operatorname{bk}(h)|$ (for every $h\in H^{\prime}$) and $\sum_{h\in H^{\prime}:\sigma^{\prime}(h)=s}|\operatorname{bk}(h)|$. It is well known that the weighted set $(S,w)$ is a good proxy for $X$, which we will use in our restricted $k$-Means subroutine.
Application: $D^{2}$-Sampling. Another more interesting application is to implement $D^{2}$-sampling, where the task is to sample an $x\in X$ with probability proportional to $\mathrm{dist}^{2}(x,S)$ [AV07]. To implement this, we compute for each high-level bucket $\operatorname{bk}(h)$ $(h\in H^{\prime})$ and the low-level buckets therein the sum of squared distance to the assigned $\sigma^{\prime}(h)$. Next, to do the sampling, first sample a high-level bucket ID $h\in H^{\prime}$ proportional to the sum-of-squared distance to $\sigma^{\prime}(h)$, then further sample a low-level bucket $\ell\in f(h)$ within $f(h)$ again proportional to the sum-of-squared distance, and finally do a uniform sample in the sampled (low-level) bucket $\operatorname{bk}(\ell)$. Here, uniform sampling works since Property 2 in Lemma 4.7 ensures points in a bucket have similar distance to $S$, and by this same property we can estimate for high-level bucket $h\in H$ the $\sum_{x\in\operatorname{bk}(h)}\mathrm{dist}^{2}(x,\sigma^{\prime}(h))$ by $|\operatorname{bk}(h)|\cdot\mathrm{dist}^{2}(x,\sigma^{\prime}(h))$ where $x$ is an arbitrary point in $\operatorname{bk}(h)$ (and note that both $|\operatorname{bk}(h)|$ and this point $x$ can be maintained), and similarly for a low-level bucket $\ell\in L$. This eventually enables us to sample $x$ proportional to $\mathrm{dist}^{2}(x,S)$, up to $O(\mathrm{poly}(\Gamma))$ factor. This approximate $D^{2}$-sampler is sufficient to be used in our another key subroutine, augmented $k$-Means.
@@
Representing Partial High-level Buckets by Low-level Buckets. The caveat is that the level-$i$ buckets $\varphi_{i}(\textsc{Ball}(s,2^{i}))$ which we add to $H^{\prime}$ should actually exclude data points already handled/deleted from levels $\leq i-1$. In other words, only a part of a high-level bucket should be added. To describe these partial bucket, we introduce low-level buckets, which are “refinement” of high-level buckets, defined as follows: for each $i$, for each bucket ID $z\in\varphi_{i}([\Delta]^{d})$, partition the level-$i$ high-level bucket $\varphi^{-1}_{i}(z)$ with respect to $\varphi_{i-1}$ (i.e., the level-$(i-1)$ high-level buckets). The set of all these parts (i.e., the intersection with $(i-1)$-level buckets) compromise the low-level buckets $L$. Now, each high-level bucket with ID $h\in H$ should maintain a list of existing/non-empty intersecting low-level buckets, in order to determine the set of points belong $h$. This is precisely our $f:H\to 2^{L}$.
Locality of Low-level Buckets. One might wonder if we need also to consider partitioning a level-$i$ high-level bucket by buckets from $i-2,i-3,\ldots$ levels and include all these into $L$ and potentially also to $f(h)$. We show that this is not necessary and only level-$(i-1)$ is needed; specifically, when we remove a high-level bucket $h$ from some level $i$, then all points in this $h$ must all be removed (possibly from levels less than $i$), without leaving partially some low-level buckets behind. This is a crucial property to control the time complexity.
@@
Time complexity. The running time of the dynamic algorithm is governed by the consistency guarantee of consistent hashing. In particular, when we examine $\varphi_{i}(\textsc{Ball}(s,2^{i}))$, we only need to consider at most $O(\Lambda)$ intersecting high-level buckets. Moreover, the consistency also guarantees that each (high-level) bucket in level-($i-1$) may result in at most $O(\Lambda)$ low-level buckets in level $i$.
@@ 4.4 Restricted and Augmented $k$-Means
We discuss how to implement the two main subroutines in the algorithm of [BCF25a]: restricted $k$-means and augmented $k$-means. The restricted $k$-means algorithm is used in Step 1, 2, 4 of the framework in Section 3.2, and augmented $k$-means is used in Step 4. Both problems take a parameter $r>0$, as well as dataset $X$ and center set $S$ as dynamic input. The restricted $k$-Means aims to remove $O(r)$ points from $S$, and augmented $k$-Means aims to add $O(r)$ points to $S$, both to minimize the resultant clustering cost.
@@ 4.4.1 Restricted $k$-Means
As a standard reduction, we may first move each point $x\in X$ to its nearest center in $S$, and then consider the problem on the induced weighted set $(S,w_{S})$ instead of the point set $X$. Therefore, in what follows, we assume $X$ is a weighted set supported on $S$.
A Sketch-Based Approach.
The high-level idea is to extract a tiny sketch of $S$, then solve the restricted $k$-Means problem on this sketch. Let $\widehat{\mathrm{dist}}$ be a (dynamic) approximate ANN distance array as in Lemma 4.6. The sketch is constructed as follows:
  1. 1.
    Pick the top $O(r)$ centers from $S$ whose $w_{S}(s)\cdot\widehat{\mathrm{dist}}^{2}(s,S-s)$ value is the smallest, denoted as $T_{1}$.
  2. 2.
    Let $T_{2}\subseteq S-T_{1}$ be a set of $O(r)$ centers such that $\forall s\in T_{1},\mathrm{dist}(s,T_{2})\leq\mathrm{poly}(\epsilon^{-1})\cdot\mathrm{dist}(s,S-T_{1})$, i.e. the set of approximate nearest neighbors in $S-T_{1}$ for all centers in $T_{1}$.
The sketch is defined as $T:=T_{1}\cup T_{2}$, and we show that $T$ approximately preserves the optimal value of the restricted $k$-means objective:
$\mathrm{OPT}^{T}_{|T|-r}(T)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathrm{poly}(\epsilon^{-1})\cdot\mathrm{OPT}^{S}_{|S|-r}(S).$
Finally, it suffices to run a vanilla $(|T|-r)$-means algorithm on $T$ and obtain a $\mathrm{poly}(\epsilon^{-1})$-approximate solution $R\subseteq T\subseteq S$:
$\textsc{cost}(R,T-R)=\textsc{cost}(T,T-R)\leq\mathrm{poly}(\epsilon^{-1})\cdot\mathrm{OPT}^{T}_{|T|-r}(T)\leq\mathrm{poly}(\epsilon^{-1})\cdot\mathrm{OPT}_{|S|-r}^{S}(S).$
This step takes $\tilde{O}(|T|^{1+\epsilon})=\tilde{O}(r^{1+\epsilon})$ time by applying static almost-linear time $k$-means algorithms [lTS24, JJLL25]. The correctness of $R$ as a solution on the full set $S$ follows from a straightforward derivation:
$\textsc{cost}(S,S-R)=\textsc{cost}(R,S-R)\leq\textsc{cost}(R,T-R)\leq\mathrm{poly}(\epsilon^{-1})\cdot\mathrm{OPT}^{S}_{|S|-r}(S).$
Dynamic Data Structure.
The key step in making the above process dynamic, is to maintain the center weights $w_{S}$. This can be solved using the cluster weight estimation structure which we introduced in Section 4.3. Once this is addressed, the construction of the sketch $T=T_{1}\cup T_{2}$ can be implemented via ANN data structures (using e.g., the one we mentioned in Section 4.2), in total time $O(r\cdot\Lambda)$.
@@ 4.4.2 Augmented $k$-Means
There exists a data structure that handles two dynamic sets $X,S\subseteq[\Delta]^{d}$, where updates are generated by an adaptive adversary. For every query with an integer $a\geq 1$, it runs in time $\tilde{O}(a\cdot 2^{\epsilon d})$ to return a set $A\subseteq[\Delta]^{d}$ of $O(a\cdot\mathrm{poly}(\epsilon^{-1})\cdot d\log\Delta)$ points such that the following holds w.h.p.:
$\textsc{cost}(X,S+A)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(1)\cdot\min_{A^{\star}\subseteq[\Delta]^{d}:|A^{\star}|=a}\textsc{cost}(X,S+A^{\star}).$
@@
Our approach is to iteratively select and add points to $S$ until a good augmentation is achieved. We borrow the idea from the classical $k$-means++ algorithm [AV07], which employs $D^{2}$-sampling – i.e., sampling a point from $X$ with probability proportional to $\mathrm{dist}^{2}(x,S)$ – to iteratively select centers and achieve an $O(\log k)$-approximation for the vanilla $k$-means problem. A follow-up work [ADK09] considers sampling $O(k)$ centers to achieve a bi-criteria $O(1)$-approximation for vanilla $k$-means, and we adapt their analysis to show that repeated $D^{2}$-sampling similarly yields a good approximate solution to the augmented $k$-means problem in our setting. We employ an approximate $D^{2}$-sampling in Section 4.3 to implement this idea in dynamic setting.

### v3

@@ 5.3 Approximate Assignment Structure
However, there is no finite ratio with small recourse for this task, since one can keep on inserting/deleting a point of $S$ to change the assignment drastically. To bypass this barrier, instead of maintaining the full assignment $\sigma:X\to S$, our idea is to hash the dataset into buckets $\varphi(X)$ and explicitly maintain only an intermediate assignment $\sigma^{\prime}$ between bucket (ID) and $S$. The final assignment $X\to S$ is implicitly defined by assigning all points in the same bucket to the same bucket partner in the intermediate assignment.
@@
Lemma 5.7 implies key lemmas that we use for restricted $k$-means and augmented $k$-means (see Section 5.4 for the definitions of these sub-tasks). Similar to Lemma 5.7, these results maintain dynamic $X$ and $S$, work against an adaptive adversary and succeed with high probability.
Application: Cluster Weight Estimation. An immediate application is to compute the injectivity $w_{S}(s)$ of points $s\in S$ in the assignment $\sigma$ (which is $\mathrm{poly}(\Gamma)$-approximate by Property 1 of Lemma 5.7), via maintaining $|\operatorname{bk}(h)|$ (for every $h\in H^{\prime}$) and $w_{S}(s):=\sum_{h\in H^{\prime}:\sigma^{\prime}(h)=s}|\operatorname{bk}(h)|$. It is well known that the weighted set $(S,w_{S})$ is a good proxy for $X$, which we will use in our restricted $k$-means subroutine.
Application: $D^{2}$-Sampling. Another more interesting application is to implement $D^{2}$-sampling, where the task is to sample an $x\in X$ with probability proportional to $\mathrm{dist}^{2}(x,S)$ [ArthurV07]. To implement this, we compute for each high-level bucket $\operatorname{bk}(h)$ $(h\in H^{\prime})$ and the low-level buckets therein the sum of squared distance to the assigned $\sigma^{\prime}(h)$. Next, to do the sampling, first sample a high-level bucket ID $h\in H^{\prime}$ proportional to the sum-of-squared distance to $\sigma^{\prime}(h)$, then further sample a low-level bucket $\ell\in f(h)$ within $f(h)$ again proportional to the sum-of-squared distance, and finally do a uniform sample in the sampled (low-level) bucket $\operatorname{bk}(\ell)$. Here, uniform sampling works since Property 2 in Lemma 5.7 ensures points in a bucket have similar distance to $S$, and by this same property we can estimate for high-level bucket $h\in H$ the $\sum_{x\in\operatorname{bk}(h)}\mathrm{dist}^{2}(x,\sigma^{\prime}(h))$ by $|\operatorname{bk}(h)|\cdot\mathrm{dist}^{2}(x,\sigma^{\prime}(h))$ where $x$ is an arbitrary point in $\operatorname{bk}(h)$ (and note that both $|\operatorname{bk}(h)|$ and this point $x$ can be maintained), and similarly for a low-level bucket $\ell\in L$. This eventually enables us to sample $x$ proportional to $\mathrm{dist}^{2}(x,S)$, up to $O(\mathrm{poly}(\Gamma))$ factor. This approximate $D^{2}$-sampler is sufficient to be used in our another key subroutine, augmented $k$-means.
@@
Representing Partial High-level Buckets by Low-level Buckets. The caveat is that the level-$i$ buckets $\varphi_{i}(\textsc{Ball}(s,2^{i}))$ which we add to $H^{\prime}$ should actually exclude data points already handled/deleted from levels $\leq i-1$. In other words, only a part of a high-level bucket should be added. To describe these partial bucket, we introduce low-level buckets, which are “refinement” of high-level buckets, defined as follows: for each $i$, for each bucket ID $z\in\varphi_{i}([\Delta]^{d})$, partition the level-$i$ high-level bucket $\varphi^{-1}_{i}(z)$ with respect to $\varphi_{i-1}$ (i.e., the level-(i1)(i-1) high-level buckets). The set of all these parts (i.e., the intersection with (i1)(i-1)-level buckets) compromise the low-level buckets $L$. Now, each high-level bucket with ID $h\in H$ should maintain a list of existing/non-empty intersecting low-level buckets, in order to determine the set of points belong $h$. This is precisely our $f:H\to 2^{L}$.
Locality of Low-level Buckets. One might wonder if we need also to consider partitioning a level-$i$ high-level bucket by buckets from $i-2,i-3,\ldots$ levels and include all these into $L$ and potentially also to $f(h)$. We show that this is not necessary and only level-(i1)(i-1) is needed; specifically, when we remove a high-level bucket $h$ from some level $i$, then all points in this $h$ must all be removed (possibly from levels less than $i$), without leaving partially some low-level buckets behind. This is a crucial property to control the time complexity.
@@
Time Complexity. The running time of the dynamic algorithm is governed by the consistency guarantee of consistent hashing. In particular, when we examine $\varphi_{i}(\textsc{Ball}(s,2^{i}))$, we only need to consider at most $O(\Lambda)$ intersecting high-level buckets. Moreover, the consistency also guarantees that each (high-level) bucket in level-(i1i-1) may result in at most $O(\Lambda)$ low-level buckets in level $i$.
@@ 5.4 Restricted and Augmented $k$-Means
We discuss how to implement the two main subroutines in the algorithm of [BCF24]: restricted $k$-means and augmented $k$-means. The restricted $k$-means algorithm is used in Step 1, 2, 4 of the framework in Section 4.2, and augmented $k$-means is used in Step 4. Both problems take a parameter $r>0$, as well as dataset $X$ and center set $S$ as dynamic input. The restricted $k$-means aims to remove $r$ points from $S$, and augmented $k$-means aims to add $r$ points to $S$, both to minimize the resultant clustering cost.
@@ 5.4.1 Restricted $k$-Means
For ease of illustration, we may assume that $X=S$; this is natural, since each point $x\in X$ can be moved to its (approximate) nearest center in $S$ with only a constant-factor increase in the clustering objective. Under this assumption, the goal of the restricted $k$-means problem becomes finding a set $R$ of $r$ points in $S$ that minimizes
$\textsc{cost}(S,S-R)\penalty 10000\ =\penalty 10000\ \textsc{cost}(R,S-R)\penalty 10000\ =\penalty 10000\ \sum_{s\in R}\mathrm{dist}^{2}(s,S-R).$
The first observation here is that the restricted $k$-means problem with $r$ centers to delete is equivalent to the $k^{\prime}$-means problem with $k^{\prime}=k-r$, where the solution is restricted to be selected from $S$. Hence, a naïve approach would be to directly apply an efficient algorithm for Euclidean $k$-means (e.g. [laTourS24, abs-2504-03513]); however, this would only achieve a superlinear $|S|^{1+\epsilon}$ running time.
Our Coreset-based Algorithm.
Our algorithm, at a high level, is to speed up this naïve approach by designing a coreset for the restricted $k$-means problem. Roughly speaking, this coreset guarantees that running an $O(1)$-approximation algorithm for the restricted $k$-means on it yields a solution whose ratio is larger but still constant for the full set $S$. We give a coreset construction with size $O(r)$, and this immediately helps to speed up the naïve approach to a nearly tight $\tilde{O}(r^{1+\epsilon})$ time while obtaining an $O(1)$-approximation for the restricted $k$-means on $S$.
This notion of coreset is new and is specifically devised for our restricted $k$-means problem. This notion is particularly different than the standard coreset for clustering in the literature (see e.g. [DBLP:conf/stoc/Har-PeledM04]), and a main conceptual difference is that ours is a coreset with respect to the points to be deleted (i.e., $R$), whereas a standard coreset preserves costs for the center set (i.e., SRS-R). Moreover, the techniques for constructing our new coreset is also very different than commonly used sampling techniques [DBLP:journals/siamcomp/Chen09, DBLP:conf/stoc/FeldmanL11, DBLP:conf/stoc/Cohen-AddadSS21, DBLP:conf/stoc/Cohen-AddadLSS22, DBLP:conf/focs/0001CPSS24, Huang0024, DBLP:conf/soda/Cohen-AddadD0SS25] as in standard coresets, and we instead employ a greedy algorithm to build the coreset.
Next, we give an overview of our coreset construction, which consists of two major steps. To start, in the first step, we greedily select $6r$ points $s\in S$ with the smallest values of $\mathrm{dist}(s,S-{s})$ as the candidate centers. Denoting the resulting set by $T_{1}$, we show (in Claim 5.9) that the candidate set $T_{1}$ contains, as a subset, a good solution to the restricted $k$-means problem.
Claim 5.9.
There exists a subset of $r$ points from $T_{1}$ that is an $O(1)$-approximation to the restricted $k$-means on the full set $S$.
However, the existence of such a good solution does not imply that $T_{1}$ is a coreset, since the clustering objective may be significantly distorted on $T_{1}$, and thus an approximation algorithm run on $T_{1}$ may not yield the good solution guaranteed by Claim 5.9. Therefore, the second step is to address this issue, and we proceed in the following way. For each selected point $s\in T_{1}$, we find its (approximate) nearest neighbors among the unselected points $S-T_{1}$. Denoting these (approximate) nearest neighbors by $T_{2}$ and letting $T:=T_{1}\cup T_{2}$, we prove the following claim.
Claim 5.10.
For any subset $R\subseteq T_{1}$, it holds that
$\textsc{cost}(S,S-R)\penalty 10000\ \leq\penalty 10000\ \textsc{cost}(T,T-R)\penalty 10000\ \leq\penalty 10000\ O(1)\cdot\textsc{cost}(S,S-R).$
Claim 5.10 implies that the objectives of all candidate solutions are preserved on $T$ within a constant factor. Therefore, by combining Claims 5.9 and 5.10, it can be shown that $T=T_{1}\cup T_{2}$ is the desired coreset.
In what follows, we explain the intuition behind Claims 5.9 and 5.10 separately, and then discuss the dynamic implementation of this construction.
Candidate Centers via Greedy Selection (Claim 5.9).
Recall that $T_{1}$ consists of the $6r$ points $s\in S$ with the smallest values of $\mathrm{dist}(s,S-{s})$. Here, we consider the case where the optimal solution $R^{*}$ to the restricted $k$-means problem on $S$ is disjoint from $T_{1}$, for ease of illustration. (The case where $R^{*}$ and $T_{1}$ overlap is handled similarly.)
The key to the proof is that, among the points in $T_{1}$ we show the existence of a subset of $r$ points, denoted by $R^{\prime}$, such that
$\forall s\in R^{\prime},\quad\mathrm{dist}(s,S-s)\penalty 10000\ \geq\penalty 10000\ \Omega(1)\cdot\mathrm{dist}(s,S-R^{\prime}).$
Since $R^{\prime}\subseteq T_{1}$ and $R^{*}\cap T_{1}=\emptyset$, and given that $T_{1}$ consists of the points $s\in S$ with the smallest $\mathrm{dist}(s,S-s)$ values, it follows that
$\forall s_{1}\in R^{\prime},\forall s_{2}\in R^{*},\quad\mathrm{dist}(s_{1},S-s_{1})\penalty 10000\ \leq\penalty 10000\ \mathrm{dist}(s_{2},S-s_{2}).$ (7)
As a result,
$\sum_{s_{2}\in R^{*}}\mathrm{dist}^{2}(s_{2},S-R^{*})\geq\sum_{s_{2}\in R^{*}}\mathrm{dist}^{2}(s_{2},S-s_{2})\geq\sum_{s_{1}\in R^{\prime}}\mathrm{dist}^{2}(s_{1},S-s_{1})\geq\Omega(1)\cdot\sum_{s_{1}\in R^{\prime}}\mathrm{dist}^{2}(s_{1},S-R^{\prime}).$
This concludes that $R^{\prime}\subseteq T_{1}$ is an $O(1)$-approximation to the restricted $k$-means problem, thereby proving Claim 5.9.
Nearest Unselected Neighbors for Candidates (Claim 5.10).
The inequality $\textsc{cost}(S,S-R)\leq\textsc{cost}(T,T-R)$ (equivalently, $\textsc{cost}(R,S-R)\leq\textsc{cost}(R,T-R)$) follows directly from the fact that $T$ is a subset of $S$. We then focus on the inequality $\textsc{cost}(T,T-R)\leq O(1)\cdot\textsc{cost}(S,S-R)$. Consider a fixed candidate solution $R^{\prime}\subseteq T_{1}$ and suppose that its restricted $k$-means objective $\textsc{cost}(T_{1},T_{1}-R^{\prime})=\textsc{cost}(R^{\prime},T_{1}-R^{\prime})$ on the candidate set $T_{1}\subseteq S$ is severely distorted, i.e., significantly larger than its objective $\textsc{cost}(S,S-R^{\prime})=\textsc{cost}(R^{\prime},S-R^{\prime})$ on the full set $S$. This distortion occurs only if, for some point $s\in R^{\prime}$, the distance $\mathrm{dist}(s,T_{1}-R^{\prime})$ is significantly larger than $\mathrm{dist}(s,S-R^{\prime})$. This further means that the nearest neighbor of $s$ in $S-R^{\prime}$ is not contained in $T_{1}-R^{\prime}$; such a neighbor must lie in $S-T_{1}$ since $R^{\prime}\subseteq T_{1}\subseteq S$. Therefore, this issue can be directly fixed by adding, for every candidate point $s\in T_{1}$, its (approximate) nearest neighbor in $S-T_{1}$ to the coreset, which is precisely our second step. This outlines the main idea behind the proof of Claim 5.10.
Dynamic Implementations.
This coreset construction is well suited for the dynamic setting, since it only relies on maintaining/querying approximate nearest neighbors, which is supported by our data structures introduced in Section 5.2. Moreover, we note that all the above arguments hold for the case where $X=S$. Therefore, we actually need an additional step that “moves” each point in $X$ to its (approximate) nearest neighbor in $S$, resulting in a weighted set $(S,w_{S})$ where the weight $w_{S}(s)$ equal to the number of points moved to $s$. To realize this moving step in the dynamic setting, we need to maintain the weight $w_{S}$, which is efficiently supported by our cluster weight estimation structure introduced in Section 5.3.
@@ 5.4.2 Augmented $k$-Means
There exists a data structure that handles two dynamic sets $X,S\subseteq[\Delta]^{d}$, where updates are generated by an adaptive adversary. For every query with an integer $r\geq 1$, it runs in time $\tilde{O}(r\cdot 2^{\epsilon d})$ to return a set $A\subseteq[\Delta]^{d}$ of $O(r\cdot\mathrm{poly}(\epsilon^{-1})\cdot d\log\Delta)$ points such that the following holds w.h.p.:
$\textsc{cost}(X,S+A)\penalty 10000\ \leq\penalty 10000\ O(1)\cdot\min_{A^{\star}\subseteq[\Delta]^{d}:|A^{\star}|=r}\textsc{cost}(X,S+A^{\star}).$
@@
Our approach is to iteratively select and add points to $S$ until a good augmentation is achieved. We borrow the idea from the classical $k$-means++ algorithm [ArthurV07], which employs $D^{2}$-sampling – i.e., sampling a point from $X$ with probability proportional to $\mathrm{dist}^{2}(x,S)$ – to iteratively select centers and achieve an $O(\log k)$-approximation for the vanilla $k$-means problem. A follow-up work [DBLP:conf/approx/AggarwalDK09] considers sampling $O(k)$ centers to achieve a bi-criteria $O(1)$-approximation for vanilla $k$-means, and we adapt their analysis to show that repeated $D^{2}$-sampling similarly yields a good approximate solution to the augmented $k$-means problem in our setting. We employ an approximate $D^{2}$-sampling in Section 5.3 to implement this idea in dynamic setting.